OpenStudy (anonymous):

given t(z) = 2z^2 -4z + 6 determine t(-6)

OpenStudy (anonymous):

just plug in -6 for z

OpenStudy (anonymous):

@cardinal_carlo okay thanks and how would I solve this one... Find the axis of symmetry of the functions graph f(x) = 2x^2 + 20x+ 46

OpenStudy (anonymous):

First, let's find the vertex. \[f'(x) = 4x + 20\]\[~~~~~~~0 = 4x + 20\]\[~ -20 = 4x\]\[~~-5 = x\] Alright, the axis of symmetry of f(x) = 2x^2 + 20x+ 46 is x = -5.

OpenStudy (anonymous):

@cardinal_carlo I got -4. Thank you so much btw